Day Shift: 06/05/99

Jason LaDue (
Sat, 5 Jun 1999 14:47:38 -0600 (MDT)

Shift Summary for Evening shift, 05/29/97

On Shift:
Jason LaDue
Valeri Jejer

Other People Present: Ed Blucher, Hogan Nguyen, Angela, Patty MacBride.

Hours of beam delivered at all intensities: 0.0
Hours of beam delivered at nominal intensity: 0
Hours of E799 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...):

Accounting of lost time:

Waiting for DC2 to be ready.

Runs ended this shift:
13373 - COSMIC run, good data
13374 - COSMIC run, Banana debugging
13375 - COSMIC run, Banana debugging
13376 - COSMIC run, Banana debugging
13377 - COSMIC run, Banana debugging
13378 - COSMIC run, Banana debugging
13379 - COSMIC run, No CsI
13380 - COSMIC run, No CsI, Special prescale for Valeri.

Tapes brought to FCC this shift: 0

Accesses during shift: Open Access

Description of Shift:

09:15 Angela reseated wire 59 in DC2. It was off by a 1000 microns on
one end. The chamber was open for about 5 minutes.
10:30 Jason L. loads in the 832 maps in the Bananas.
11:50 ktev2 crashes. The crash writes out a dump that fills up the
the root partion. That file was deleted.
12:00 Ed swaps HV divider for channels 2344 and 2712. Also swaps the
dpmt for channel 2344.
12:45 Angela and Brian put in He Bag #5 and notice that it is about
2 feet too short.
13:15 Notice that the humidity in the Hall is at about 60%, Call
the on-duty mechanic and he resets the Munters.
14:30 V0 RF_L1 timing fixed so that most hits are in Slice 5.
15:00 John Peoples visits.
15:24 Notice that Temp has gone red on the SDAQ screen, power is off.
Notifying Experts.
15:30 Ed and Hogan are still downstairs in the Block House stringing
a fiber.

run spill time comment
13373 07:45 Ended run
13379 10:08 Started run
13379 11:50 ktev2 crashes, ending the run.
13380 14:00 Started run